OpenStudy (anonymous):

solve and graph the compound inequality 7x + 8 > or = to - 48 or 5x + 2 > or = to - 8 the solution of the compound inequality is {x|x>=to ---?

OpenStudy (anonymous):

solve and graph the compound inequality 7x + 8 > or = to - 48 or 5x + 2 > or = to - 8 the solution of the compound inequality is {x|x>=to ---?